OpenStudy (junyang96):

is there any way to integrate sec^3 without using integration by parts?

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\sec^3 x~dx}\) \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{\cos x}{(1-\sin^2 x)^2}~dx}\) u= sin x du = cos(x) dx \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{1}{(1-u^2)^2}~du}\) http://www.wolframalpha.com/input/?i=partial+fractions+1%2F%28%281-u%5E2%29%5E2%29 then continue on from partial fraction decomposition (that I show in wolfram)